You have just turned 40 and you plan to save for retirement. You plan to retire...

  1. You have just turned 40 and you plan to save for retirement. You plan to retire in 20 years. Once you retire you would like to have an income of $80,000 per year for the next 10 years. Determine the amount you must deposit at the beginning of each year to finance your retirement income. Use the following assumptions to determine this annual deposit:

  1. All savings compound at a rate of 12% per year.
  2. You make the first deposit today (when you turn 40) and the last deposit on the day you turn 59 (20 payments).
  3. You make the first withdrawal when you turn 60 and the last withdrawal when you turn 69 (10 withdrawals).
Answer #1
PVAnnuity Due = c*((1-(1+ i)^(-n))/i)*(1 + i )
C = Cash flow per period
i = interest rate
n = number of payments
PV= 80000*((1-(1+ 12/100)^-10)/(12/100))*(1+12/100)
PV = 506259.98
FVAnnuity Due = c*(((1+ i)^n - 1)/i)*(1 + i )
C = Cash flow per period
i = interest rate
n = number of payments
506259.98= Cash Flow*(((1+ 12/100)^20-1)/(12/100))*(1+12/100)
Cash Flow = 6273.46